On aircraft carriers, catapults are used to accelerate jet aircraft to flight speeds in a short distance. One such catapult takes a 18,500-kg jet from 0 to 68 m/s in 2.9 s. (Assume the catapult acts in the positive horizontal direction. Indicate the direction with the sign of your answer where appropriate.)

(a) What is the acceleration of the jet in m/s2? 23.45 m/s2

What is the acceleration of the jet in g's? g

(b) How far does the jet travel while it is accelerating? m

(c) How large is the force that the catapult must exert on the jet? N
Answer #1

18 S00 k (a) me V- 6 8 S V= u+a 68ms0 a (29s) a- 23.4STmls2 acceleation in d a23.4ml x9. 2-392 X u 늘지2345m)e) x(2.9s) 2 X = O48 SOo k68ls -0) 2. S F 433793 N